OpenStudy (anonymous):

You have 57.0 mL of a 0.400 M stock solution that must be diluted to 0.100 M. Assuming the volumes are additive, how much water should you add?

OpenStudy (nechirwan):

\[M _{1}\times V _{1} = M _{1}\times V _{1}\] \[4 \times 57=1\times V _{2}\] \[V _{2} =228 ml\] \[V _{water added} =228-57=171ml\]